G varies inversely with the square of H. When H is 4, G is 0.25. Fing the value of G when H is 0.3
Answer by ewatrrr(24785)

G = k/H^2
.25 = k/16
.25*16 = k
4 = k
G = 4/H^2 when h = .3
G = 4/.09 = 400/9 = 44 + 4/9